OpenStudy (anonymous):

5/9v+w=2 solve for v

OpenStudy (jack1):

2 variables, one equation, you can only solve for v in terms of w

OpenStudy (anonymous):

from my son's 9th grade math book

OpenStudy (anonymous):

\[5\frac{9}{v}+w=2\] or\[5\frac{9}{v+w}=2\] which one is the problem

OpenStudy (anonymous):

top

OpenStudy (anonymous):

solving for v means you want v on one side... since lets get rid of the w on the left side since we want v on the left side. subtract w to both sides

OpenStudy (anonymous):

nope... 5/9 v+w=2

OpenStudy (anonymous):

yes -w on the left

OpenStudy (anonymous):

\[ \frac 5 9 v+w = 2 \]

OpenStudy (anonymous):

5/(9(2-w))=v

OpenStudy (anonymous):

Don't use this site to cheat. Your son needs to do this problem.

OpenStudy (anonymous):

1st subtract w so you get 5/9v = 2-w 2nd multiply v on both sides getting 5/9 = v(2-w) 3rd divide both sides by 2-w getting 5/(18-9w) = v

OpenStudy (anonymous):

Thanks Propulsion

OpenStudy (anonymous):

Don't cheat. There is no point. If the kid doesn't get it, get them a tutor.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."